A soap bubble film (n=1.42) of thickness 106 nm re
Answer #1

Constructive interference occurs here when -

2 tc = \lambdan / 2, 3 (\lambdan / 2), 5 (\lambdan / 2)

we know that, tc = (\lambdan / 4) = 106 nm

The next three thicknesses of this film which will be given as :

tc' = 3 (106 nm) = 318 nm

tc'' = 5 (106 nm) = 530 nm

tc''' = 7 (106 nm) = 742 nm